The following is an example of a linear graph.
A scuba diver is 400 ft. below sea level and rising at the rate of 4 ft every 3 seconds.

Write an equation to represent the depth of the diver (letting y=0 represent the water surface).

Can any one help me to translate that in to equation..
Thanks

Answer by nerdybill(7384)

A scuba diver is 400 ft. below sea level and rising at the rate of 4 ft every 3 seconds. Write an equation to represent the depth of the diver (letting y=0 represent the water surface).
.
"Slope-intercept" form of a line:
y = mx + b
where
m is the slope
b is the y-intercept
.
From:"A scuba diver is 400 ft."
We know that at time (t) equal 0, depth is -400 feet (below sea level) -- this is your y-intercept.
.
From:"rising at the rate of 4 ft every 3 seconds"
We get our slope (m) as 4/3
.
Therefore, our equation is
d(t) = (4/3)t + (-400)
d(t) = (4/3)t - 400
where
d(t) is the depth in feet
t is time in secs